At a given instant an 59 mH inductor has a 40 mA. What is the energy stored in the inductor at this instant?

________
Answer #1

A)
U = 0.5*L*I^2

= 0.5*59*10^-3*(40*10^-3)^2

= 47.2 micro J

B)

I should be sqrt(27) times


let T is the time taken.

sqrt(27)*40*10^-3 = 40*10^-3 + T*90*10^-3


T = (sqrt(27)*40 - 40)/90

= 1.865 s